OpenStudy (anonymous):

Factor completely x2 – x – 6

OpenStudy (swissgirl):

(x-3)(x+2)

OpenStudy (swissgirl):

2*-3=-6 2-3=-1

OpenStudy (anonymous):

x2 – x – 6=0 X(X+2)-3(X+2)=0 (X-3)(X+2)=0 X-3=0 or X+2=0 X=3 or X=-2